# Estimating Body and Hand Motion in an Ego-sensed World

### v1

@@ 1 Introduction
Head-mounted devices are permeating the mainstream market, with applications in areas like virtual reality, content creation, and assistive technologies. For research in machine perception, these devices also present exciting opportunities in the form of egocentric sensing. Sensors from wearable devices provide abundant observations of 3D environments, while capturing the embodied perspective of human agents as they observe, navigate, and interact with the world around them.
@@
We therefore introduce EgoAllo, a system that uses egocentric inputs to estimate the wearer’s actions in the allocentric world. This is a difficult task: while body parts like hands occasionally appear in egocentric image frames, most body parameters are never directly observed by these devices. Accurate estimates in the global frame also require harmony between both pose and height parameters. These must be consistent with both the egomotion and scene scale, aligning the feet to the ground and the head to the sensed height of the camera.
We cast estimation as guided sampling from a diffusion-based motion model. We consider two egocentric inputs: a proprioceptive one in the form of head pose, which we use for conditioning, and images, which we use for visual hand observations that are incorporated via guidance. We then estimate body pose, height, and hand parameters, unlike most prior works in egocentric human motion estimation [47, 6] that focus on body pose.
We achieve this by training a head motion-conditioned diffusion model as a motion prior, and guiding its sampling to recover a hand-body sequence that aligns with image observations. Our results are enabled by a key insight: the head observation conditioning representation is critical for accurate ego-sensed pose estimation. We study the role of choices for this representation by (1) identifying desirable spatial and temporal invariance properties that are not fulfilled by existing systems, (2) using these properties to derive improved parameterizations for our motion prior, and (3) presenting a systematic evaluation that shows between 4.9% and 17.9% error difference when compared against prior work. Furthermore, we show how the resulting system can improve hand estimation, reducing errors by over 40% compared to noisy monocular estimates.
@@ 2 Related Work
3D human recovery from external visual inputs. A large body of work has addressed estimating the parameters of human body models like SCAPE [2] or SMPL and its variants  [51, 77, 61] from third-person visual inputs, where human subjects are observed from the view of outside cameras. The majority of these works focus on extracting 3D representations from single images, for example by lifting 2D keypoint observations to 3D [56], via end-to-end regression [33, 39, 58, 20, 75, 31, 60], via optimization [61, 19, 43], or by exploiting synergies between regression and optimization [41]. When multiple frames are available in the form of a video, temporal context and tracking can also be incorporated [69, 16, 34, 106, 38, 62, 63]. The inputs (images) and outputs (human meshes) of many of these systems are superficially similar to the egocentric setting addressed by EgoAllo, but egocentric devices present unique challenges because the body being estimated is typically behind the outwards-facing cameras used as input.
@@
Denoising diffusion for human motion. The core of EgoAllo is a denoising diffusion model [82, 23, 67] from which we can sample 4D human body motion. While diffusion models are primarily known for their success in text-conditioned image generation [76, 78], they have also enabled advances in human motion synthesis conditioned on modalities like text [36, 109, 35], music [91, 1], poses [47, 35], and object geometry [42, 46, 48]. EgoAllo adopts a similar conditional diffusion approach, while specifically studying the design of conditioning parameters used for ego-sensed human motion estimation. The iterative nature of denoising diffusion also enables guidance [13, 85, 35, 111, 30, 11], where denoising steps are steered to satisfy a desired objective. We use guidance to incorporate observations like visual hand pose observations during test-time.
@@
Human motion from egocentric observations. EgoAllo builds on intuition from several prior works in egocentric sensing for human motion estimation. Many rely on fisheye cameras that place the wearer’s body into the field of view [27, 90, 89, 93, 96, 74, 89, 94]. Other approaches rely on body-mounted cameras [81], simulation-based physical plausibility [104, 105, 53], body- and hand-mounted inertial sensors [102, 103, 45], handheld controllers [6, 28, 29], and interaction cues from other humans [57]. Concurrent works have also used the Nymeria [54] dataset for egocentric motion with language description outputs [24], as well for online settings with scene geometry and CLIP [68] feature inputs [21]. Most relevantly, EgoEgo [47] demonstrates how human body poses can be estimated offline without body observability assumptions. The authors accomplish this by carefully integrating several components: a monocular SLAM system [87], a pose-conditioned gravity vector regression network, an optical flow feature-conditioned head orientation and scale regression network, and a head pose-conditioned body diffusion model. For estimates that can be more easily grounded in the allocentric coordinate frame, EgoAllo differs in both inputs—we study conditioning parameters computed from the metric SLAM poses provided by devices like Project Aria [83]—and outputs—we consider body height variation and hand poses.
@@
Conditioning for ego-sensed poses. Prior works vary in how head pose information is parameterized and used as neural network input. AvatarPoser [28] and BoDiffusion [6] parameterize head pose as four components: world-frame orientation, orientation deltas, world-frame position, and world-frame position deltas. These works are focused on settings with VR controller input, and parameterize controller pose inputs the same way. EgoEgo [47]’s diffusion model uses only absolute head positions and orientations, but, similar to HuMoR [72], in implementation defines a per-sequence canonical coordinate frame to ensure that all input trajectories passed to the model are aligned with the same initial $xy$ position and forward direction. In our work, we will refer to this as sequence canonicalization. Finally, EgoPoser [29] proposes a similar scheme that aligns initial positions for both head pose and controller pose inputs. We propose an alternative to these parameterizations that is motivated by the robustness and generalization benefits of invariance, as observed in prior work for designing both representations [95, 52, 79, 9, 112, 101] and neural network architectures [44, 32, 107, 88, 8, 7, 12, 14, 98]. Specifically, we introduce in Section 3.1.2 a parameterization that is invariant to both spatial and temporal shifts.
@@ 3 Method
We study the problem of using sensors from an egocentric device to estimate the actions of a wearer in an allocentric coordinate frame. We assume a flat floor and two egocentric inputs—poses from the device’s SLAM system and camera images.
@@ 3.1.2 Invariant conditioning
Invariance goals. As discussed in Section 2, prior work varies in how the function $g$ is implemented. To understand how they impact the learning problem, we propose two invariance properties for head motion representations. Each reduces representational redundancy, which eases the learning problem.
@@
Head motion representations for a given body motion should be independent of location within a temporal window. This can be expressed as temporal shift equivariance. Let $\vec{c}^{\ t}$ be as defined in Equation 2. If we shift inputs by $\delta$ such that $\{\vec{c}^{\ 1}_{\text{shift}},\dots,\vec{c}^{\ T}_{\text{shift}}\}=g(\{\mathbf{T}_{\text{world},\text{cpf}}^{1+\delta},\dots,\mathbf{T}_{\text{world},\text{cpf}}^{T+\delta}\})$, then $g$ should satisfy $\vec{c}^{\ t}_{\text{shift}}=\vec{c}^{\ t+\delta}$ for overlapping timesteps.
@@
Relative transforms alone, however, do not encode information relative to the scene or floor: full trajectories can even be flipped upside down without impacting $\Delta\mathbf{T}_{\text{cpf}}^{t-1,t}$. We therefore propose to ground relative motion to the floor plane with a transformation between the CPF frame and a per-timestep canonical frame, which is computed by projecting the CPF frame to the floor. This encodes head height and orientation:
@@ 3.2 Estimation via sampling
where a transformer [92] $\mu_{\theta}$ is trained to predict the posterior mean from $\vec{x}_{n}$ and conditioning $\vec{c}$. With noise-dependent weight term $w_{n}$, the loss can be written as:
$\min_{\hat{x}_{0}}\mathbb{E}_{\vec{x}_{0}\sim q_{0}}\mathbb{E}_{n\sim\mathcal{U},\vec{x}_{n}\sim q_{n|0}}\left[w_{n}\lVert\mu_{\theta}(\vec{x}_{n},n,\vec{c})-\vec{x}_{0}\rVert^{2}\right].$ (9)
After training, we estimate human motions by following DDIM [84] for conditional sampling. The final EgoAllo sampling procedure includes several additional components: a global alignment phase, guidance losses for physical constraints and visual hand observations, and a path fusion [3] approach for longer sequence lengths. We describe these below.
@@ 3.2.2 Guidance losses
We begin by running HaMeR on the egocentric image corresponding to each timestep $t$. When detected, this produces 3D hand estimates in the form of MANO [77] joint parameters and camera-centric 3D hand keypoints $\hat{\mathbf{t}}^{t}_{\text{camera,j}}$ for hand joint set $j\in\mathcal{H}$. With each subcripted $\lambda$ indicating a scalar weighting term, we have:
@@
The 3D objective $\mathcal{E}_{\text{hands3D}}^{(\Theta)}$ minimizes the distance between the detected MANO hand parameters and the corresponding SMPL-H hand parameters, in terms of wrist pose and local joint rotations. With $\Pi_{K}$ as projection with camera intrinsics $K$, $\mathbf{p}_{world,j}^{(\Theta^{t})}\in\mathbb{R}^{3}$ as the world position for joint $j$ at time $t$, and $\mathbf{T}_{\text{camera},\text{cpf}}$ from the device calibration, the reprojection loss is:
@@ 4 Experiments
Training. To train EgoAllo models used in our experiments, we need sequences containing human body and hand pose parameters, body shapes, and device SLAM poses $\textbf{T}_{\text{world},\text{cpf}}^{t}$. Similar to prior work [47, 6, 28], we train EgoAllo using AMASS [55] with synthesized device poses. We annotate train split sequences by anchoring a central pupil frame between vertices corresponding to the left and right pupils in the blend skinned mesh, and at train time sample sequences uniformly between length 32 and 128.
@@ 4.1 Body estimation
In our first set of experiments, we evaluate body estimation from only device SLAM poses, without considering images or hands. This setting allows us to isolate the advantages of our body motion prior, while facilitating direct comparison against methods that do not consider hands [47].
@@ 4.1.1 Invariant conditioning evaluation
We train conditional diffusion models with otherwise identical architecture using each parameterization, and then evaluate on the AMASS [55] test set. Metrics, including percent difference compared to EgoAllo, are reported in Table 1.
Overall, we find that the choice of conditioning parameterization makes a dramatic impact on estimation accuracy. We observe accuracy improve consistently as invariance properties are incorporated into the representation. Compared to EgoAllo, Absolute conditioning would increase MPJPE by over 23% for both shorter (length 32) and longer (length 128) sequences. Compared to EgoAllo, SeqCanonical conditioning would increase MPJPE by nearly 18% for length 32 sequences and 12% for length 128 sequences.
@@ 4.1.2 Comparisons against baselines
Shape estimation evaluation. To better understand the shape estimation characteristics of EgoAllo, we compare against against the “mean” shape used by EgoEgo and the NoShape ablation. On the AMASS test set, we find: EgoAllo slightly improves overall shape ($19\text{mm}\to 18\text{mm}$ mean vertex-to-vertex error) and produces much better height ($52\text{mm}\to 32\text{mm}$ mean height error), but is not able to generalize in terms of body weight ($5\text{kg}\to 8\text{kg}$ mean weight error). The body shape is inferred from the wearer’s head pose, which intuitively provides strong height constraints but is less corelated with weight. Accurate height is key for proper scene placement, as reflected by both the MPJPE and GND metrics.
@@ 5 Dicussion
Limitations and future work. While the core contributions of EgoAllo are general, the current implementation of our system has a few limitations that we hope to explore in future work. First, diffusion model guidance is a test-time optimization process that requires hyperparameters and incurs a runtime cost: our optimization uses a CPU-based sparse linear solver [10], which for length-128 sequences takes several seconds for each guidance step. In the future, it may be possible to bootstrap using outputs from our model to train a feedforward model that avoids this step. Success for hand guidance also still depends on reasonable monocular hand estimates. Estimation can therefore fail as a result of errors like left/right flipping or spurious detections, which we found causes high errors in our Table 3 hand estimation metrics. Finally, we also train only on AMASS [55], which includes floor planes but no detailed scene geometry. As a result, our method will fail if we cannot detect an approximate floor plane; this is most likely in settings like hills or staircases. In the future, we hope to extend our insights to data with more detailed scene information, which concurrent work has highlighted the usefulness of in informing human body estimation [21].
@@ A.1 Floor height estimation
A key requirement of our algorithm is SLAM poses that can be situated relative to the floor. While floor heights are provided in our training data, they are not directly available on real-world data. We found that a simple RANSAC-based algorithm works well on real-world data from Project Aria [59]. We filter SLAM points by confidence, then use RANSAC to find a z-value with that best fits a plane. Example floor plane outputs, using scenes from the EgoExo4D [17] dataset, are shown in Figure 7.
@@ A.2.1 Biomech57 evaluation details
The one exception is the Aria Digital Twins dataset [59], which we use for quantitative body metrics. Each device wearer in the Aria Digital Twins dataset is recorded via an Optitrack motion capture system, which records 57 joint locations (30 hand joints, 27 body joints) following the Biomech57 joint template. The Biomech57 joint standard does not match the SMPL-H joints we use in our method. To evaluate our method on ADT, we match and compare the common major joints between the two templates.
We manually corresponded each of the 57 joints between Biomech57 and the standard SMPL-H joint conventions. While the majority of these have 1:1 correspondences—feet, knees, hips, shoulders, elbows, wrist, and finger joints, for example, are consistently defined—others like the head and collar bone joints are misaligned. When computing MPJPE values for the ADT dataset, we mask out these joints.

### v2

@@ 1 Introduction
Head-mounted devices are permeating the mainstream market, with applications in areas like virtual reality, content creation, and assistive technologies. For research in machine perception, these devices also present exciting opportunities in the form of egocentric sensing. Sensors from wearable devices provide abundant observations of 3D environments, while capturing the embodied perspective of human agents as they navigate and interact with the world around them.
@@
We therefore introduce EgoAllo, a system that uses egocentric inputs to estimate the wearer’s actions in the allocentric frame of the world. This is a difficult task: while body parts like hands occasionally appear in egocentric image frames, most body parameters are never directly observed by these devices. Accurate estimates in the global frame also require harmony between both pose and height parameters. These must be consistent with the egomotion and scene scale, aligning estimated feet to the ground and head to the sensed camera height.
We cast estimation as guided sampling from a human motion diffusion model. We consider two inputs: head poses from SLAM and visual hand observations from images. We then estimate body pose, height, and hand parameters, unlike most prior works in egocentric human motion estimation [46, 6] that focus on body pose.
We achieve this by training a head motion-conditioned diffusion model as a motion prior, and guiding its sampling to recover a hand-body sequence that aligns with image observations. Our results are enabled by a key insight: that the representation used for head pose conditioning is critical for accurate ego-sensed motion estimation. We study choices for this representation by (1) identifying desirable spatial and temporal invariance properties that are not fulfilled by existing systems, (2) using these properties to derive improved parameterizations for our motion prior, and (3) presenting a systematic evaluation that shows between 4.9% and 17.9% error difference when compared against prior work. Furthermore, we show how the resulting system can improve hand estimation, reducing errors by over 40% compared to noisy monocular estimates.
@@ 2 Related Work
3D human recovery from external visual inputs. A large body of work has addressed estimating the parameters of human body models like SCAPE [2] or SMPL and its variants  [50, 76, 60] from third-person visual inputs, where human subjects are observed from the view of outside cameras. The majority of these works focus on extracting 3D representations from single images, for example by lifting 2D keypoint observations to 3D [55], via end-to-end regression [32, 38, 57, 19, 74, 30, 59], via optimization [60, 18, 42], or by exploiting synergies between regression and optimization [40]. When multiple frames are available in the form of a video, temporal context and tracking can also be incorporated [68, 15, 33, 106, 37, 61, 62]. The inputs (images) and outputs (human meshes) of many of these systems are similar to the egocentric setting addressed by EgoAllo, but egocentric devices present unique challenges because the body being estimated is typically behind the outwards-facing cameras used as input.
@@
Denoising diffusion for human motion. The core of EgoAllo is a denoising diffusion model [81, 22, 66] from which we can sample 3D human body motion. While diffusion models are primarily known for their success in text-conditioned image generation [75, 77], they have also enabled advances in human motion synthesis conditioned on modalities like text [35, 110, 34], music [90, 1], poses [46, 34], and object geometry [41, 45, 47]. EgoAllo adopts a similar conditional diffusion approach, while specifically studying the design of conditioning parameters used for ego-sensed human motion estimation. The iterative nature of denoising diffusion also enables guidance [12, 84, 34, 112, 29, 10], where denoising steps are steered to satisfy a desired objective. We use guidance to incorporate observations like visual hand pose observations during test-time.
@@
Human motion from egocentric observations. EgoAllo builds on intuition from several prior works in egocentric sensing for human motion estimation. Many rely on fisheye cameras that place the wearer’s body into the field of view [26, 89, 88, 92, 95, 73, 88, 93]. Other approaches rely on body-mounted cameras [80], simulation-based physical plausibility [104, 105, 52], body- and hand-mounted inertial sensors [102, 103, 44], handheld controllers [6, 27, 28], and interaction cues from other humans [56]. Concurrent works have also used the Nymeria [53] dataset for egocentric motion with language description outputs [23], as well for online settings with scene geometry and CLIP [67] feature inputs [20]. Most relevantly, EgoEgo [46] demonstrates how human body poses can be estimated offline without body observability assumptions. The authors accomplish this by carefully integrating several components: a monocular SLAM system [86], a pose-conditioned gravity vector regression network, an optical flow feature-conditioned head orientation and scale regression network, and a head pose-conditioned body diffusion model. EgoAllo differs in both inputs—we study conditioning parameters computed from the metric SLAM poses provided by devices like Project Aria [82]—and outputs—we consider body height variation and hand poses.
@@
Conditioning for ego-sensed poses. Prior works vary in how head pose information is parameterized and used as neural network input. AvatarPoser [27] and BoDiffusion [6] parameterize head pose as four components: world-frame orientation, orientation deltas, world-frame position, and world-frame position deltas. These works are focused on settings with VR controller input, and parameterize controller pose inputs the same way. EgoEgo [46]’s diffusion model uses only absolute head positions and orientations, but, similar to HuMoR [71], in implementation defines a per-sequence canonical coordinate frame to ensure that all input trajectories passed to the model are aligned with the same initial $xy$ position and forward direction. In our work, we will refer to this as sequence canonicalization. Finally, EgoPoser [28] proposes a similar scheme that aligns initial positions for both head pose and controller pose inputs. We propose an alternative to these parameterizations that is motivated by the robustness and generalization benefits of invariance, as observed in prior work for designing both representations [94, 51, 78, 9, 113, 101, 108] and neural network architectures [43, 31, 107, 87, 8, 7, 11, 13, 97, 98]. Specifically, we introduce in Section 3.1.2 a parameterization that is invariant to both spatial and temporal shifts.
@@ 3 Method
We study the problem of using sensors from an egocentric device to estimate the actions of the wearer in an allocentric coordinate frame. We assume a flat floor and two egocentric inputs: poses from the device’s SLAM system and camera images.
@@ 3.1.2 Invariant conditioning
Invariance goals. As discussed in Section 2, prior work varies in how the function $g$ is implemented. To understand how choices impact learning, we propose two invariance properties for head motion representations. Each reduces representational redundancy, which eases the learning problem.
@@
Head motion representations for a given body motion should be independent of location within a temporal window. This can be expressed as temporal shift equivariance. Let $\vec{c}^{\ t}$ be as defined in Equation 2. For any shift $\delta$ such that $\{\vec{c}^{\ 1}_{\text{shift}},\dots,\vec{c}^{\ T}_{\text{shift}}\}=g(\{\mathbf{T}_{\text{world},\text{cpf}}^{1+\delta},\dots,\mathbf{T}_{\text{world},\text{cpf}}^{T+\delta}\})$, $g$ should satisfy $\vec{c}^{\ t}_{\text{shift}}=\vec{c}^{\ t+\delta}$ for overlapping timesteps.
@@
Relative transforms alone, however, do not encode information relative to the scene or floor: full trajectories can even be flipped upside down without impacting $\Delta\mathbf{T}_{\text{cpf}}^{t-1,t}$. We therefore propose to ground relative motion to the floor plane with a transformation between the CPF frame and a per-timestep canonical frame, which is computed by projecting the CPF frame to the floor. This encodes head height and orientation. Our full representation then becomes:
@@ 3.2 Estimation via sampling
where a transformer [91] $\mu_{\theta}$ is trained to predict the posterior mean from noised sample $\vec{x}_{n}$ and conditioning $\vec{c}$. With noise-dependent weight term $w_{n}$, the loss can be written as:
$\min_{\theta}\ \ \mathbb{E}_{\vec{x}_{0}}\mathbb{E}_{n\sim\mathcal{U}}\left[w_{n}\lVert\mu_{\theta}(\vec{x}_{n},n,\vec{c})-\vec{x}_{0}\rVert^{2}\right].$ (9)
After training, we estimate human motions by following DDIM [83] for sampling. The final EgoAllo sampling procedure includes several additional components: a global alignment phase, guidance losses for physical constraints and visual hand observations, and a path fusion [3] approach for longer sequence lengths. We describe these below.
@@ 3.2.2 Guidance losses
We begin by running HaMeR on the egocentric image corresponding to each timestep $t$. When detected, this produces 3D hand estimates in the form of MANO [76] joint parameters and camera-centric 3D hand keypoints $\hat{\mathbf{p}}^{t}_{\text{camera,j}}$ for hand joint set $j\in\mathcal{H}$. Optionally, wrist and palm poses can also be estimated using Project Aria’s Machine Perception Services [82]. With each subcripted $\lambda$ indicating a scalar weighting term, we have:
@@
The 3D objective $\mathcal{E}_{\text{hands3D}}^{(\Theta)}$ minimizes the distance between the detected hand parameters and the corresponding SMPL-H hand parameters, in terms of wrist pose and local joint rotations. With $\Pi_{K}$ as projection with camera intrinsics $K$, $\mathbf{p}_{world,j}^{(\Theta^{t})}\in\mathbb{R}^{3}$ as the world position for joint $j$ at time $t$, and $\mathbf{T}_{\text{camera},\text{cpf}}$ from the device calibration, the reprojection loss is:
@@ 4 Experiments
Training. To train EgoAllo models used in our experiments, we need sequences containing human body and hand pose parameters, body shapes, and device SLAM poses $\textbf{T}_{\text{world},\text{cpf}}^{t}$. Similar to prior work [46, 6, 27], we train EgoAllo using AMASS [54] with synthesized device poses. We annotate train split sequences by anchoring a central pupil frame between vertices corresponding to the left and right pupils in the blend skinned mesh, and at train time sample sequences between length 32 and 128.
@@ 4.1 Body estimation
In our first set of experiments, we evaluate body estimation from only device SLAM poses, without considering images or hands. This setting allows us to isolate the advantages of our body motion prior, while facilitating direct comparison against methods that do not consider hands.
@@ 4.1.1 Invariant conditioning evaluation
We train conditional diffusion models with otherwise identical architecture using each parameterization, and then evaluate on the AMASS [54] test set. Metrics and percent differences compared to EgoAllo are reported in Table 1.
Overall, we find that the choice of conditioning parameterization makes a dramatic impact on estimation accuracy. We observe accuracy improve consistently as invariance properties are incorporated into the representation. Compared to EgoAllo, Absolute conditioning increases MPJPE by over 23% for both shorter (length 32) and longer (length 128) sequences. Compared to EgoAllo, SeqCanonical conditioning increases MPJPE by nearly 18% for length 32 sequences and 12% for length 128 sequences.
@@ 4.1.2 Comparisons against baselines
Shape estimation evaluation. To better understand the shape estimation characteristics of EgoAllo, we compare against against the “mean” shape used by EgoEgo and the NoShape ablation. On the AMASS test set, we find: EgoAllo slightly improves overall shape ($19\text{mm}\to 18\text{mm}$ mean vertex-to-vertex error) and produces much better height ($52\text{mm}\to 32\text{mm}$ mean height error), but is not able to generalize in terms of body weight ($5\text{kg}\to 8\text{kg}$ mean weight error). The body shape is inferred from the wearer’s head pose, which intuitively provides strong height constraints but is less correlated with weight. Accurate height is key for proper scene placement, as reflected by both the MPJPE and GND metrics.
@@ 5 Dicussion
Limitations and future work. While the core contributions of EgoAllo are general, the current implementation of our system has a few limitations that we hope to explore in future work. First, diffusion model guidance is a test-time optimization process that depends on hyperparameters and incurs a runtime cost. In the future, it may be possible to bootstrap using outputs from our model to train a feedforward model that avoids this step. Success for hand guidance also still depends on reasonable monocular hand estimates. Estimation can therefore fail as a result of errors like left/right flipping or spurious detections, which we found causes high errors in our Table 3 hand estimation metrics. Finally, we also train only on AMASS [54], which includes floor planes but no detailed scene geometry. As a result, our method will fail if we cannot detect an approximate floor plane; this is most likely in settings like hills or staircases. In the future, we hope to extend our insights to data with more detailed scene information, which concurrent work has highlighted the usefulness of in informing human body estimation [20].
@@ A.1 Floor height estimation
A key requirement of our algorithm is SLAM poses that can be situated relative to the floor. While floor heights are provided in our training data, they are not directly available on real-world data. We found that a simple RANSAC-based algorithm works well on real-world data from Project Aria [58]. We filter SLAM points by confidence, then use RANSAC to find a z-value with that best fits a plane. Example floor plane outputs using scenes from the EgoExo4D [16] dataset are shown in Figure 7.
@@ A.2.1 Biomech57 evaluation details
The one exception is the Aria Digital Twins dataset [58], which we use for quantitative body metrics. Each device wearer in the Aria Digital Twins dataset is recorded via an Optitrack motion capture system, which records 57 joint locations (30 hand joints, 27 body joints) following the Biomech57 joint template. To evaluate our method on ADT, we match and compare the common major joints between the two templates. We manually corresponded each of the 57 joints between Biomech57 and the standard SMPL-H joint conventions. While the majority of these have 1:1 correspondences—feet, knees, hips, shoulders, elbows, wrist, and finger joints, for example, are consistently defined—we mask out others like the head and collar bone joints that are misaligned.